OpenStudy (anonymous):

csc2 θ = 8 cot θ + 10

OpenStudy (anonymous):

can anyone please help

OpenStudy (anonymous):

can you show me the steps to how to solve this?

OpenStudy (shubhamsrg):

you should note that csc^2 x = 1 + cot^2 x after you re arrange some terms , you'll see a quadratic equation formed in cot x

hartnn (hartnn):

yes, is that \(\csc^2 \theta =8 \cot \theta+10\) ?? and you need to solve for thetha ? any interval given ?

OpenStudy (anonymous):

\[\csc ^{2}\theta=8\cot \theta+10\] \[1+\cot ^{2}\theta=8\cot \theta+10\] \[\cot ^{2}\theta-8\cot \theta-9=0\] 1*-9=-9 9-1=8 \[\cot ^{2}\theta-\left( 9-1 \right)\cot \theta-9=0\] now you can solve.